# Raising the Bar of AI-generated Image Detection with CLIP

### v1

@@ 4.4 Beyond low-level forensic traces
Figure 5: Top (from left to right): a high resolution synthetic image generated by Stable Diffusion XL [58] and its 4$\times$ decimated version; a real image and the corresponding image obtained at the output of Stable Diffusion [62] used as an autoencoder. Bottom: Fourier spectra of the noise residuals for images shown on the top. A suitable decimation removes Fourier peaks in synthetic images, while passing a real image through an autoencoder creates new peaks.
GAN family Diffusion family AVG
Pro
Style
Style
Style
Giga
Score
ADM GLIDE
Latent
Stable
DeepFl.
Ediff-I DiT SDXL
Method
real/synth. training
size (k)
GAN
GAN2
GAN3
GANT
GAN
SDE
Diff.
Diff.
IF
Wang et al.
LSUN / ProGAN
360 / 360
100.
96.5
89.0
98.9
66.6
32.9
64.3
48.5
59.2
41.5
78.0
64.9
58.6
54.3
68.1
PatchFor.
CelebA,FF / various
84 / 272
92.3
84.5
87.9
91.2
64.7
83.3
74.8
96.2
78.1
62.4
62.7
78.7
83.1
68.4
79.2
Grag. et al.
LSUN / ProGAN
360 / 360
100.
99.8
98.3
98.8
82.8
92.1
74.7
62.8
91.9
52.5
69.9
69.6
65.3
58.0
79.8
Liu et al.
LSUN / ProGAN
360 / 360
100. 99.8
97.8
98.5
98.2 95.4
82.5
76.5
97.6
77.4
72.2
98.7
88.0
31.1
86.7
Corvi et al.
COCO,LSUN / Latent
180 / 180
79.4
73.7
38.4
97.1
39.9
65.0
80.7
91.9
100. 99.7 99.9
85.7
100. 100.
82.2
LGrad
LSUN / ProGAN
72 / 72
100.
91.2
87.3
81.8
82.2
80.6
76.9
66.1
81.1
61.5
68.8
74.1
56.2
57.2
76.1
Ojha et al.
LSUN / ProGAN
360 / 360
100.
93.9
90.9
98.2
96.0
58.4
86.7
80.8
85.7
89.5
92.9
80.6
77.8
85.1
86.9
DIRE-1
LSUN-Bed / ADM
40 / 40
50.6
56.9
49.3
99.9
74.1
44.3
75.7
71.4
68.7
39.4
98.9
99.1
99.6
47.1
69.7
DIRE-2
LSUN-Bed / St.GAN
40 / 40
54.2
52.5
44.3
99.6
76.0
41.0
70.1
70.1
69.3
46.9
97.0
98.2
98.3
42.8
68.6
Ours 1k
COCO / Latent
1 / 1
98.7
89.8
91.9
100.
81.2
89.4
80.3
99.9 94.0
87.4
96.8
98.7
93.9
89.2
92.2
Ours 1k+
COCO / Latent
1 / 1
89.2
79.7
83.2
99.7
74.2
84.0
74.9
99.6
82.0
89.5
97.8
99.2
91.8
90.3
88.2
Ours 10k
COCO / Latent
10 / 10
99.8
92.5
92.7
100.
84.4
88.0
83.2
99.9
96.2
91.8
97.6
99.2
95.2
88.8
93.5
Ours 10k+
COCO / Latent
10 / 10
87.1
81.8
81.9
99.8
73.5
85.7
79.1
99.3
82.5
90.6
98.4
99.7
93.2
90.8
88.8
Table 3: Comparison with SOTA methods in terms of AUC. For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+). Results obtained on the very same dataset used for training are in light gray, while entries in bold underline the best OOD performance for each dataset. We also report the mean AUC by averaging the AUC scores over all datasets.
GAN family Diffusion family AVG
Pro
Style
Style
Style
Giga
Score
ADM GLIDE
Latent
Stable
DeepFl.
Ediff-I DiT SDXL
Method
real/synth. training
size (k)
GAN
GAN2
GAN3
GANT
GAN
SDE
Diff.
Diff.
IF
Wang et al.
LSUN / ProGAN
360 / 360
100.
69.2
56.2
71.5
52.8
50.4
50.3
50.7
51.3
50.1
50.0
50.4
50.1
49.8
57.3
PatchFor.
CelebA,FF / various
84 / 272
64.8
50.9
50.1
51.1
51.3
50.0
50.3
50.5
51.3
50.2
50.0
50.0
51.2
50.0
51.5
Grag. et al.
LSUN / ProGAN
360 / 360
100. 96.6 92.6
91.9
69.8
83.8
54.5
56.8
67.5
48.9
48.9
52.1
49.2
51.4
68.9
Liu et al.
LSUN / ProGAN
360 / 360
100.
96.2
78.6
89.4
93.3 89.2
67.2
73.2
84.3
65.3
50.5
92.0
64.3
39.6
77.4
Corvi et al.
COCO,LSUN / Latent
180 / 180
50.0
50.7
50.0
52.0
57.7
50.0
54.5
57.3
99.8 99.7
64.3
50.9
99.8 99.7
66.9
LGrad
LSUN / ProGAN
72 / 72
100.
77.2
62.3
68.5
74.7
60.3
61.2
55.4
69.3
61.5
60.9
62.6
53.2
58.7
66.1
Ojha et al.
LSUN / ProGAN
360 / 360
99.9
77.3
82.4
87.1
81.2
57.1
60.9
57.1
68.1
72.6
65.7
66.0
55.4
68.0
71.3
DIRE-1
LSUN-Bed / ADM
40 / 40
51.4
50.8
50.0
96.5
73.1
50.0
73.2
73.2
73.3
49.7
95.3 95.6
94.9
49.9
69.8
DIRE-2
LSUN-Bed / St.GAN
40 / 40
51.4
50.7
50.0
82.0
66.6
50.0
65.9
65.9
66.4
49.6
80.7
81.5
82.2
49.6
63.7
Ours 1k
COCO / Latent
1 / 1
69.7
72.9
82.0
98.9
69.0
76.6
71.2
97.8 86.9
70.9
63.2
80.3
81.0
60.2
77.2
Ours 1k+
COCO / Latent
1 / 1
68.4
70.3
73.5
96.2
62.9
72.8
65.8
94.7
74.5
77.1
81.6
91.9
78.0
75.2
77.4
Ours 10k
COCO / Latent
10 / 10
90.1
83.0
84.0
98.7
66.5
75.6
72.3
93.7
89.9
72.9
59.5
78.8
76.8
54.4
78.3
Ours 10k+
COCO / Latent
10 / 10
73.8
72.3
71.4
95.0
60.8
75.6
68.2
90.5
74.4
77.3
81.7
92.4
77.4
70.0
77.2
Table 4: Comparison with SOTA methods in terms of Accuracy, fixed 0.5 threshold. Results obtained on the very same dataset used for training are in light gray, while entries in bold underline the best OOD performance for each dataset. We also report the mean Accuracy by averaging the Accuracy scores over all datasets.
5 Comparison with the state-of-the-art
In this Section we perform an extensive comparison with SoTA methods in different scenarios. For our approach we show four variants: trained on 1k real and 1k fake images; 10k real and 10k fake images; trained on 1k and 1k fake images but including compressed/resized images (1k+) and trained on 10k and 10k fake images but including compressed/resized images (10k+). In our analysis we only include SoTA techniques with code and/or pre-trained models publicly available on-line in order to ensure a fair comparison. We considered the following approaches:
@@ 5 Comparison with the state-of-the-art
    Grag et al. [33] proposes a simple modification to the ResNet50 architecture which allows to better preserve low-level forensic traces and is trained on the same dataset introduced in [71].
@@
  7. (g)
    Ojha et al. [56] uses CLIP-based features and perfoms detection using the CLIP feature space. Both nearest neighbor and linear probing are tested, even if the latter gives better performance.
@@
    DIRE [72] is based on the idea that synthetic images can be reconstructed better than real images by a pre-trained model. To this end it is proposed to train a ResNet-50 in two different ways, using ADM images (DIRE 1) and StyleGAN images (DIRE 2), respectively.
Comparison over GAN- and DM-based generators. In Tab. 3 and Tab. 4 we show results in terms of AUC and Accuracy, respectively, over 14 GAN-based and DM-based generators proposed in the scientific literature and perfectly known. For each method, results obtained on the very same dataset used for training are in light gray, to signal an ID scenario, of little interest for our analysis. Entries in bold, instead, underline the best OOD performance for each dataset. Focusing on AUC data, Tab. 3, we observe that methods trained on a GAN dataset work generally well on other dataset of the same family, and much worse on DM datasets. Then, the roles switch for methods trained on DM datasets. This is nothing new, as generators of the two families share architectural details that leave similar traces on the generated images. No SoTA method works uniformly well on all datasets. On the contrary, the proposed lightweight CLIP-based detector provides a uniformly good performance, and the version with 10k+10k reference images without augmentation outperforms the best competitor (notably, again a CLIP-based method) by +6.6%, obtaining an average AUC well above 93.5% a truly impressive accomplishment.
In terms of Accuracy, Tab. 4, performance lowers significantly, and in some cases dramatically. The proposed CLIP-based detectors keep providing the best performance, generally much better than SoTA methods, except for Liu’s method that performs almost at the same level.
Comparison over commercial tools. However, the performance of Liu’s method, as well as that of several other methods, drops catastrophically when considering the Commercial Tools (DALL·E2, DALL·E3, Midjourney, FireFly) both in terms oa AUC and Accuracy, see Tab. 5. Needless to say, this is the most realistic and interesting scenario, often with images of unknown origin and no prior information on the possible generation process. In this situation, most methods provide unreliable decisions. Corvi’s method works surprisingly well on some datasets, maybe generated by diffusion models similar to Latent. However, only the proposed CLIP-based approach provides a stable good performance in all cases. Note that, now, the versions with resized and recompressed images in the reference set perform significantly better than those without augmentation.
Robustness to perturbations. The trend observed over unknown models is further accentuated when images are subject to some post-processing, Tab. 6. These impairments tend to wash out forensic traces, to the point that most SoTA methods become basically useless with coin-tossing performance, especially on the unknown commercial models. Again, the CLIP-based detectors provide a more than decent performance, further supporting our conjecture on the high-level nature of the forensic traces they rely upon.
Commercial tools
DALL·E2
DALL·E3
Midjourney
FireFly
Average
Method
AUC/Acc
AUC/Acc
AUC/Acc
AUC/Acc
AUC/Acc
Wang et al.
64.8 / 50.2
10.9 / 49.4
40.2 / 49.5
84.8 / 50.3
50.2 / 49.8
PatchFor.
41.9 / 50.0
52.7 / 50.0
57.8 / 50.0
49.4 / 50.0
50.5 / 50.0
Grag. et al.
58.3 / 49.2
2.4 / 36.9
43.1 / 40.5
63.5 / 47.5
41.8 / 43.6
Liu et al.
70.4 / 66.4
0.2 / 30.8
40.7 / 43.3
11.8 / 31.0
30.8 / 42.9
Corvi et al.
69.4 / 50.1
60.8 / 50.0
100. / 100.
98.0 / 51.1
82.1 / 62.8
LGrad
58.6 / 55.5
37.9 / 47.8
56.3 / 57.9
40.6 / 47.5
48.3 / 52.1
Ojha et al.
95.2 / 83.5
36.4 / 47.3
66.2 / 52.4
97.5 / 89.9
73.8 / 68.3
DIRE-1
44.7 / 50.0
47.6 / 50.0
51.0 / 49.7
57.4 / 49.9
50.2 / 49.9
DIRE-2
41.0 / 49.9
49.6 / 50.0
47.8 / 49.8
43.0 / 49.8
45.3 / 49.9
Ours 1k
89.1 / 62.2
77.6 / 53.4
74.3 / 53.2
77.1 / 52.1
79.5 / 55.2
Ours 1k+
83.9 / 68.2
94.8 / 83.1
79.7 / 64.7
86.0 / 68.4
86.1 / 71.1
Ours 10k
88.8 / 57.8
82.9 / 51.1
75.4 / 51.2
79.9 / 50.9
81.7 / 52.8
Ours 10k+
84.5 / 63.9
94.4 / 76.6
83.8 / 62.7
87.6 / 64.5
87.6 / 67.0
Table 5: AUC/Accuracy on Commercial Tools. We show AUC and Accuracy on very recent synthetic generators: DALL·E2, DALL·E3, Midjourney v5 and Firefly.
Families of Generators
GAN
Diffusion
Comm. Tools
Average
Method
AUC/Acc
AUC/Acc
AUC/Acc
AUC/Acc
Wang et al.
78.9 / 62.1
59.3 / 50.4
44.1 / 49.9
60.8 / 54.1
PatchFor.
54.3 / 50.0
64.8 / 50.3
62.9 / 50.1
60.7 / 50.1
Grag. et al.
89.2 / 68.5
67.9 / 50.8
47.3 / 50.0
68.1 / 56.4
Liu et al.
54.7 / 51.6
56.3 / 51.1
53.0 / 50.7
54.6 / 51.1
Corvi et al.
66.7 / 54.1
76.1 / 62.1
52.1 / 50.1
65.0 / 55.4
LGrad
51.2 / 51.0
48.0 / 50.5
49.4 / 50.4
49.5 / 50.7
Ojha et al.
86.9 / 74.9
73.7 / 59.8
49.6 / 50.6
70.1 / 61.8
DIRE-1
48.9 / 50.3
51.0 / 50.5
50.1 / 49.9
50.0 / 50.2
DIRE-2
49.1 / 50.0
52.3 / 52.5
45.3 / 49.9
48.9 / 50.8
Ours 1k
76.4 / 68.7
81.2 / 71.4
74.3 / 61.9
77.3 / 67.3
Ours 1k+
75.3 / 67.8
87.8 / 77.6
84.2 / 66.0
82.4 / 70.4
Ours 10k
76.8 / 60.7
82.7 / 59.6
73.8 / 50.5
77.8 / 56.9
Ours 10k+
78.2 / 67.9
88.7 / 77.0
83.0 / 65.6
83.3 / 70.2
Table 6: AUC/Accuracy in the presence of post-processing. All the images have been randomly cropped, resized and compressed to simulate a realistic scenario over the web.
6 Discussion
In this work, we have shown a series of good features of CLIP that can help detecting real from synthetic content. Through an extensive experimental analysis we found that:
  • Maximizing the diversity of reference CLIP features impacts positively on performance, even when a relatively large number of examples is considered.
  • It is well known that CLIP descriptive power descends from its huge pre-training set. Further increasing this set keeps boosting the performance, up to $10\%$.
  • Experimental evidence seems to show that CLIP features, even when tuned to forensic applications, are mostly independent from low-level forensic traces. This opens the way to interesting developments. On one hand, one can gain immunity to possible malicious attacks working on low-level artifacts. On the other hand, there is room for further performance improvement by suitable fusion with classic detectors.
Overall, the present study provides a number of interesting insights but leaves much room for future work. We believe that new and better forensic methods can be proposed on top of CLIP. One major area of study can be to develop few-shot methods that adapt the detector on the fly to the situation of interest. Future work should also consider interpretability, to begin with understanding which forensic features is the detector exploiting to make its decisions.
Acknowledgment.
We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, a TUM-IAS Rudolf Mößbauer Fellowship and a Google Gift. This material is also based on research sponsored by the Defense Advanced Research Projects Agency (DARPA) and the Air Force Research Laboratory (AFRL) under agreement number FA8750-20-2-1004. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright notation thereon. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of DARPA or the U.S. Government. In addition, this work has received funding by the European Union under the Horizon Europe vera.ai project, Grant Agreement number 101070093. It is also supported by the PREMIER project, funded by the Italian Ministry of Education, University, and Research within the PRIN 2017 program. Finally, we want to thank Koki Nagano for useful discussions, and Tero Karras, Yogesh Balaji, Ming-Yu Liu for sharing data for StyleGAN-T and eDiff-I experiments.
Supplemental Material
In this appendix, we report additional ablation studies (Sec. A) and robustness analysis (Sec. B). We also enlarge our initial dataset with additional synthetic generators and carry out further experiments on generalization (Sec. C). Furthermore, we conduct more experiments on the ability of our approach and SoTA methods to deal with possible attacks from malicious users (Sec. D). Finally, we show some experiments carried out on social networks in a few-shot scenario (Sec. E).
Figure 6: Average AUC for our approach considering different classifiers and different features on the original dataset. Top: 1k images, Bottom: 10k images.
Figure 7: Average AUC for our approach considering different classifiers and different features on the dataset after random compression and resizing. Top: 1k images, Bottom: 10k images.
Appendix A Additional ablation study
In Section 4 of the paper we describe the CLIP-based detection procedure used in all experiments. Two important design choices outlined in the paper are that i) we extract features from the next to last layer of the CLIP ViT L/14 architecture; and ii) we use a SVM classifier, designed based on a limited set of reference features. Here we test some alternative solutions, that is:
@@ Appendix A Additional ablation study
In Figures 6 and 7, we show the results in terms of average AUC over the dataset described in Section 3 (main paper) using both original images and images impaired by common post-processed steps such as recompression and resizing. For all detectors, we consider the versions 1k and 10k (1000 and 10000 reference images, respectively).

### v2

@@ 6 Beyond low-level forensic traces
Figure 5: Top (from left to right): a synthetic image generated by Stable Diffusion XL [59] and its 4$\times$ decimated version; a real image and the corresponding image processed by the autoencoder of Stable Diffusion XL. Bottom: Fourier spectra of the noise residuals for images shown on the top. A suitable decimation removes Fourier peaks in synthetic images, while passing a real image through an autoencoder creates new peaks.
Refer to captionRefer to caption
Figure 6: Scatter plots of scores provided by Corvi et al. ($y$-axis) and proposed method without and with augmentation ($x$-axis) on 2400+2400 images sampled from the 18 datasets of Tab.1.
7 Discussion
We have proposed a simple method based on CLIP features to distinguish real images from synthetic images. Through extensive experimental analysis, we found that:
  • Maximizing the diversity of reference CLIP features has a positive impact on performance, even when a relatively large number of examples is considered.
  • It is known that CLIP’s descriptive power comes from its huge pre-training set. Further increasing this set keeps boosting performance, up to $10\%$.
  • Experimental evidence suggests that CLIP features, even when adapted to forensic applications, are largely independent of low-level forensic traces. This provides some immunity to malicious attacks targeting low-level artifacts and paves the way for further performance improvements through suitable fusion with traditional detectors.
The present study provides a number of interesting insights but leaves much room for future work. We believe that new and better forensic methods can be proposed based on CLIP features. One major area of study is the development of few-shot methods that adapt the detector to the situation of interest on the fly. Future work should also consider interpretability, starting with understanding which forensic features the detector exploits to make its decisions.
downsample by 1/4 add low-level traces
Method St.Diff. St.Diff. SDXL St.Diff. St.Diff. SDXL St.Diff. St.Diff. SDXL
1.4 2.0 1.4 2.0 1.4 2.0
Corvi et al. 100. 100. 100. 41.6 46.2 45.6 79.3 80.4 67.6
Ours 1k 88.8 87.0 87.8 85.8 82.2 79.3 89.3 82.5 80.8
Ours 1k+ 90.2 89.4 88.9 89.8 91.1 90.0 90.8 86.1 83.2
Ours 10k 93.8 90.6 87.4 86.8 84.4 80.5 93.7 86.2 80.6
Ours 10k+ 94.0 90.4 90.1 95.9 92.6 91.6 94.1 86.7 84.0
Table 6: Results in terms of AUC without attacks (first group of columns) and after attacks on synthetic images (second group) and real images (third group). Corvi et al., based on low-level forensic traces, is severely affected by the attacks while the proposed method keeps working well in all conditions.
Families of Generators
GAN Diffusion Comm. Tools Average
Method AUC/Acc AUC/Acc AUC/Acc AUC/Acc
Corvi et al. 72.7 / 52.1 91.5 / 75.1 82.1 / 62.8 82.1 / 63.3
Ours 10k 92.4 / 79.1 92.6 / 73.3 80.5 / 52.6 88.5 / 68.3
Ours 10k+ 89.3 / 74.9 91.8 / 77.2 87.0 / 67.3 89.4 / 73.1
Ours fusion 92.9 / 80.1 96.9 / 87.8 88.2 / 65.3 92.7 / 77.8
Ours fusion+ 89.9 / 75.5 96.8 / 88.8 92.3 / 77.2 93.0 / 80.5
Corvi et al. 74.0 / 55.1 77.3 / 62.1 52.1 / 50.1 67.8 / 55.8
Ours 10k 76.9 / 63.6 81.1 / 63.7 73.0 / 52.4 77.0 / 59.9
Ours 10k+ 78.2 / 69.0 89.3 / 78.7 84.7 / 66.4 84.1 / 71.4
Ours fusion 78.5 / 66.8 85.1 / 71.4 72.7 / 52.6 78.7 / 63.6
Ours fusion+ 79.6 / 70.8 91.7 / 80.7 84.5 / 66.6 85.3 / 72.7
Table 7: AUC/Accuracy results for Corvi et al., proposed method (10k and 10k+ versions), and their fusion over the three families of generators: GAN, Diffusion, Commercial Tools. Top: original images; bottom: compressed/resized images.
Acknowledgment.
We gratefully acknowledge the support of this research by a TUM-IAS Hans Fischer Senior Fellowship, a TUM-IAS Rudolf Mößbauer Fellowship and a Google Gift. This material is also based on research sponsored by the Defense Advanced Research Projects Agency (DARPA) and the Air Force Research Laboratory (AFRL) under agreement number FA8750-20-2-1004. In addition, this work has received funding by the European Union under the Horizon Europe vera.ai project, Grant Agreement number 101070093. Finally, we want to thank Koki Nagano, Tero Karras, Yogesh Balaji and Ming-Yu Liu for sharing data for StyleGAN-T and eDiff-I experiments.
Supplemental Material
In this appendix, we provide a brief description of the methods used for comparison (Sec. A), report additional ablation studies (Sec. B), additional results (Sec. C), and robustness analysis (Sec. D). We also show some experiments carried out on social networks in a few-shot scenario (Sec. E). Finally, we enlarge our initial dataset with additional synthetic generators and carry out further experiments on generalization (Sec. F).
Appendix A Reference methods
In our experimental analysis we included the following methods:
@@ Appendix A Reference methods
    Gragnaniello et al. [34] proposes a simple modification to the ResNet50 architecture which allows to better preserve low-level forensic traces and is trained on the same dataset introduced in [73].
@@
    Mandelli et al. [49] relies on the ensemble of five EfficientNet-B4 networks trained on different datasets. At test-time the scores of randomly selected patches are aggregated: if at least one patch is detected as synthetic, then the entire image is classified as synthetic.
  4. (d)
@@
    Ojha et al. [57] uses a large dataset of real and synthetic images to train a simple classifier working on pre-trained CLIP features.
  9. (i)
@@
    DIRE [74] is based on the idea that synthetic images can be reconstructed better than real images by a pre-trained model. To this end a ResNet-50 is trained in two different ways, using ADM images (DIRE 1) and StyleGAN images (DIRE 2), respectively.
  10. (j)
    NPR [71] works on residual images computed as the difference between the original image and its interpolated version. The classifier is a ResNet-50 trained on only 4 classes of the ProGAN dataset.
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 7: Average AUC for our approach considering different classifiers and different features on the original dataset (top) and after random compression and resizing (bottom). Left: 1k images, Right: 10k images.
Appendix B Additional ablation study
In Section 4 of the paper we describe the CLIP-based detector used in all experiments. In the proposed procedure two key design choices are made: i) we extract features from the next to last layer of the CLIP ViT L/14 architecture; and ii) we use a SVM classifier, trained on a limited set of reference features. Here we test alternative solutions, that is:
@@ Appendix B Additional ablation study
In Figure 7, we show the results in terms of average AUC over the dataset described in Section 3 (main paper) using both original images and images impaired by common post-processing steps such as recompression and resizing. For all detectors, we consider the versions 1k and 10k (1000 and 10000 reference images per class, respectively).
@@
Families of Generators
GAN Diffusion Comm. Tools Average
Method AUC/Acc AUC/Acc AUC/Acc AUC/Acc
Wang et al. 92.1 / 69.4 55.8 / 50.3 50.2 / 49.8 66.0 / 56.5
PatchFor. 84.9 / 53.6 76.4 / 50.4 50.5 / 50.0 70.6 / 51.3
Grag. et al. 95.8 / 90.1 70.8 / 57.0 41.8 / 43.6 69.5 / 63.6
Mand. et al. 88.9 / 81.6 55.8 / 53.7 22.6 / 32.9 55.7 / 56.1
Liu et al. 99.0 / 89.7 79.9 / 69.5 30.8 / 42.9 69.9 / 67.4
Corvi et al. 72.7 / 52.1 91.5 / 75.1 82.1 / 62.8 82.1 / 63.3
LGrad 87.8 / 76.4 69.2 / 60.4 48.3 / 52.1 68.4 / 63.0
Ojha et al. 96.1 / 85.4 82.0 / 63.4 73.8 / 68.3 84.0 / 72.4
DIRE-1 65.9 / 64.4 71.6 / 72.8 50.2 / 49.9 62.5 / 62.4
DIRE-2 65.1 / 60.1 70.4 / 65.8 45.3 / 49.9 60.3 / 58.6
NPR 89.5 / 79.7 82.0 / 68.1 49.3 / 50.1 73.6 / 66.0
Ours 1k 91.2 / 76.4 92.1 / 76.2 76.6 / 54.5 86.6 / 69.0
Ours 1k+ 86.2 / 74.2 89.9 / 79.9 85.3 / 72.6 87.1 / 75.6
Ours 10k 92.4 / 79.1 92.6 / 73.3 80.5 / 52.6 88.5 / 68.3
Ours 10k+ 89.3 / 74.9 91.8 / 77.2 87.0 / 67.3 89.4 / 73.1
Table 8: Comparison with SOTA methods in terms of AUC and Accuracy. We report the mean AUC and Accuracy for each family of generators and the global average.
Families of Generators
GAN Diffusion Comm. Tools Average
Method AUC/Acc AUC/Acc AUC/Acc AUC/Acc
Wang et al. 79.2 / 62.0 59.3 / 50.4 44.1 / 49.9 60.9 / 54.1
PatchFor. 54.2 / 50.0 64.8 / 50.3 62.9 / 50.1 60.7 / 50.1
Grag. et al. 89.0 / 67.6 67.9 / 50.8 47.3 / 50.0 68.1 / 56.1
Mand. et al. 69.2 / 55.1 50.6 / 50.8 44.7 / 49.6 54.9 / 51.9
Liu et al. 54.4 / 51.5 56.3 / 51.1 53.0 / 50.7 54.6 / 51.1
Corvi et al. 74.0 / 55.1 77.3 / 62.1 52.1 / 50.1 67.8 / 55.8
LGrad 52.0 / 50.9 48.0 / 50.5 49.4 / 50.4 49.8 / 50.6
Ojha et al. 86.1 / 73.7 73.7 / 59.8 49.6 / 50.6 69.8 / 61.4
DIRE-1 47.6 / 50.2 51.0 / 50.5 50.1 / 49.9 49.6 / 50.2
DIRE-2 47.7 / 49.6 52.3 / 52.5 45.3 / 49.9 48.4 / 50.7
NPR 48.7 / 50.2 50.0 / 49.8 56.0 / 50.1 51.6 / 50.0
Ours 1k 76.2 / 68.9 80.6 / 71.2 72.2 / 60.7 76.3 / 66.9
Ours 1k+ 75.0 / 67.2 87.8 / 77.8 83.4 / 65.7 82.1 / 70.3
Ours 10k 76.9 / 63.6 81.1 / 63.7 73.0 / 52.4 77.0 / 59.9
Ours 10k+ 78.2 / 69.0 89.3 / 78.7 84.7 / 66.4 84.1 / 71.4
Table 9: AUC/Accuracy in the presence of post-processing. All images have been randomly cropped, resized and compressed to simulate a realistic scenario on the web.
Appendix C AUC vs Accuracy
In this Section we present additional results on the synthetic generators analyzed in the main paper aggregated by family (GAN, Diffusion, Commercial tools). Results are not only in terms of AUC but also of balanced accuracy computed with a fixed threshold of 0.5. Indeed in a practical scenario a threshold must be set to discriminate between real and fake images, which is not a trivial task as can be seen in Table 8. In fact, in the absence of an adequate calibration procedure, a significant drop between AUC and accuracy is often observed. However, our approach provides good results even in this challenging context, especially on post-processed images (Tab. 9) where most methods are equivalent to flipping coins.
Refer to captionRefer to captionRefer to caption